# Efficient Time Series Processing for Transformers and State-Space Models through Token Merging

### v1

@@ A.2 Experiments
Here we list additional information concerning our experimental settings and resources.
@@
Datasets We base our experiments on commonly used multivariate time series datasets covering different forecasting applications: ETTh1 and ETTm1 consist of variates measuring the power load and temperature of electric transformers in hourly and quarter-hourly granularity [11].
@@
Weather consists of meteorological quantities such as air temperature and is recorded every minutes in .111https://www.bgc-jena.mpg.de/wetter/
@@
Electricity measures the energy demand of consumers every hour [6].
@@
Traffic consists of sensors in the San Francisco Bay Area measuring the road occupancy hourly [6].
We use the same data splits for training, validation and test as [13] for consistency.
@@
Since the Chronos foundation model operates univariately and requires considerable computational resources, we randomly sample the same time series from the test set for all Chronos evaluations. For the ETTh1 dataset we do not observe relevant differences when comparing the results to the full test set.
@@
To explore token merging in an additional sequence-based domain and on a second task, we use the Dummy Mouse Enhancers Ensembl dataset [37] for classifying genomic data. It contains very long sequences of nucleotides from a mouse.
@@
Hyperparameter optimization For each transformer architecture, model size, and dataset we train models without token merging doing hyperparameter tuning of learning rate and dropout using HEBO [45]. Here we apply token merging during inference-time only. We choose the best model based on its validation MSE. We train models with the found hyperparameters, the minimum possible , and different uniformly spaced until all tokens are merged. We again choose the best model based on the MSE for further evaluation. We do hyperparameter optimization trials of both chosen models, trained with and without token merging, using HEBO to find token merging inference hyperparameters and on the validation set. Finally, we evaluate once on the test set to report our results.
@@
Hyperparameters In table 4 we list the most relevant hyperparameters we used for training the transformer models including the vanilla Transformer, Autoformer, FEDformer, Informer and Non-stationary Transformer. For training and testing HyenaDNA [9] and for testing Chronos [7] we used their default hyperparameters.
@@
Token merging in pretrained models | A6000 | 80 |
@@
Token merging during training | A6000 | 50 |
@@
Token merging improves model performance | A6000 | 30 |
@@
Dynamic merging | A6000 | 140 |
Token merging in state-space models | A6000 | 40 |
A100 | 6 |
A.3 Scaling to large models
In this section we show complete results on applying token merging to Chronos, a time series foundation model.
A.4 When does token merging improve model performance
We find token merging to have a smoothing effect improving MSE and show our results on all datasets here.
A.5 Dependencies on input length
Here we show an additional evaluation on applying token merging in Chronos models with different input lengths.

### v3

@@ Appendix C Experiments
Here, we list additional information concerning our experimental settings and resources.
@@
Datasets We base our experiments on commonly used multivariate time series datasets covering different forecasting applications: ETTh1 and ETTm1 consist of variates measuring the power load and temperature of electric transformers in hourly and quarter-hourly granularity (Zhou et al., 2021).
@@
Weather consists of meteorological quantities, such as air temperature, and is recorded every minutes in .111https://www.bgc-jena.mpg.de/wetter/
@@
Electricity measures the energy demand of consumers every hour (Godahewa et al., 2021).
@@
Traffic consists of sensors in the San Francisco Bay Area measuring the road occupancy hourly (Godahewa et al., 2021).
We use the same data splits for training, validation, and test as Wu et al. (2021) for consistency.
@@
Since the Chronos foundation model operates univariately and requires considerable computational resources, we randomly sample the same time series from the test set for all Chronos evaluations. For the ETTh1 dataset, we do not observe relevant differences when comparing the results to the full test set in figure 9.
@@
To explore token merging in an additional sequence-based domain and on a second task, we use the Dummy Mouse Enhancers Ensembl dataset (Grešová et al., 2023) for classifying genomic data. It contains very long sequences of nucleotides from a mouse.
Applying token merging In our experiments, we
generally find it beneficial to allow self-attention to transfer information between tokens before merging them. Therefore, we apply token merging between self-attention and the MLP in all transformer encoders as Bolya et al. (2023). Many transformers exhibit quadratic attention, imposing considerable computational cost. As a result, we do not find the token merging algorithm to introduce a substantial additional overhead. Thus, we choose to profit from a global merging pool for transformer encoders. For our main experiments, we also apply our causal local merging with in the transformer decoders between self-attention and cross-attention and finally unmerge all decoder tokens.
Therefore, we utilize different merging strategies in transformer encoders and decoders. In architectures utilizing additional tensors like attention masks or positional biases, we merge them using the same correspondences.
In state-space models, we merge tokens after the Hyena or Mamba operator and choose to not introduce an operation with quadratic complexity into the architecture.
@@
Hyperparameter optimization For each transformer architecture, model size, and dataset, we train models without token merging doing hyperparameter tuning of learning rate and dropout using HEBO (Cowen-Rivers et al., 2022). Here, we apply token merging during inference-time only. We choose the best model based on its validation MSE. We train models with the found hyperparameters, the minimum possible , and different uniformly spaced until all tokens are merged. We again choose the best model based on the MSE for further evaluation. We do hyperparameter optimization trials of both chosen models, trained with and without token merging, using HEBO to find token merging inference hyperparameters and on the validation set. Please note that and might be different for local merging in the encoder and causal local merging in the decoder. Finally, we evaluate once on the test set to report our results.
@@
Hyperparameters In table 6 we list the most relevant hyperparameters we used for training the transformer models, including the vanilla Transformer, Autoformer, FEDformer, Informer, and Non-stationary Transformer. For training and testing HyenaDNA (Nguyen et al., 2023) and for testing Chronos (Ansari et al., 2024), we used their default hyperparameters.
Reproducibility of measurements We report all results on the same Nvidia A6000 GPU. For training, we utilize Nvidia V100 and A100 GPUs. We measure the end-to-end inference time of the models using warm-ups and measurement runs per batch. The standard deviation of the inference time is generally in our experiments. Besides the inference time as practically most relevant quantity, we report FLOPs as a more hardware-independent measure using the thop library (Zhu, 2022). We choose the maximum possible batch size and standardize the results.
@@
| Local merging in pretrained models | A6000 | 100 |
@@
| Local merging during training | A6000 | 50 |
@@
| Token merging in state-space models | A6000 | 40 |
| A100 | 12 | |
| Dynamic token merging | A6000 | 140 |
| Improvement of forecasting quality | A6000 | 30 |
@@
Appendix D Results
Here, we show additional results for our main experiments.
D.1 Scaling to large models
In this section, we show complete results on applying token merging to Chronos, a time series foundation model.
Appendix E Further investigations
Here, we show additional experiments and results.
E.1 Token similarity measures
Different distance measures can be utilized to determine similar tokens for merging. Here, we explore the and norms as magnitude-aware metrics and the cosine similarity measuring the angular distance. Our results show that the cosine similarity outperforms both the and norms marginally. Bolya et al. (2023) further ablate the similarity metric for the vision domain.
E.2 Token pruning
Here, we compare merging tokens to pruning tokens. Generally, pruning is associated with a higher loss of information. This is also evident in our results in figure 16, where local merging outperforms local pruning.
E.3 Tokenization methods
Tokenization involves splitting the input time series into smaller units and embedding them in high-dimensional space. In our experiments, we utilize varying token representations: Transformer architectures in section 5.1 leverage multivariate tokens and a continuous embedding space, while Chronos utilizes univariate tokens embedded into discrete space. Autoformer and FEDformer transform tokens to the frequency domain. In table 8, we further include PatchTST (Nie et al., 2023), which embeds fixed-length subsequences as tokens. While most architectures generate tokens from time series, Hyena and Mamba tokenize nucleotide sequences. We argue that the tokenization method is of minor importance for token merging. Throughout all of our experiments, local merging consistently performs well on top of all token types.
| Dataset | Layers | PatchTST | ||
| MSE | Accel. | MSEΔ | ||
| ETTh1 | 2 | |||
| ETTm1 | 2 | |||
| Weather | 2 | |||
E.4 Improvement of forecasting quality — selective low-pass filter
We find token merging to have a smoothing effect, improving MSE, and show our results on all datasets here.
E.5 Improvement of forecasting quality — dataset properties
In the following, we show the frequency spectrum of ETTh1, ETTm1, Weather, Electricity, and Traffic datasets.
E.6 Redundancy of input tokens
First, we investigate whether scaling the number of input tokens increases average redundancy on the ETTh1 dataset. As demonstrated in figure 19, the same relative number of tokens is merged for a given merging threshold, independent of input length. Therefore, we suggest scaling the number of merged tokens in each layer linearly with the input length.
Positional embeddings add information about the location of a token within a sequence. As a result, two identical tokens without positional embeddings may show considerable differences when positional embeddings are included, potentially preventing merging. However, figure 19 shows that this effect on token merging is only marginal.
It is worth noting that the attention of the transformer acts as a high-dimensional low-pass filter, effectively generating more redundancy throughout the transformer layers, as Marin et al. (2021) show. Therefore, token merging not only relies on redundancy in the input data but also exploits redundancy that is introduced by the transformer itself.
E.7 Dependencies on input length
Here we show an additional evaluation on applying token merging in Chronos models with different input lengths. It is beneficial to choose a larger input length with token merging over a smaller one without.